OpenStudy (anonymous):

Use substitution to solve the system of equations. y=x-19 2x-7=27

OpenStudy (anonymous):

You look at the first equation and see that there are two variables. Then, you look at the second equation and see that there is only one variable. Because there is only one variable, you can safely assume that you can actually get a numeric value for x, whereas having 2 variables results in you have another variable in your definition of a variable. That said, You want to simply the second equation first.

OpenStudy (anonymous):

2x-7=27 +7 +7 ---------- 2x=34 x=17

OpenStudy (anonymous):

If x=17, you can substitute it in the first equation to get y=17-19. Solve that and you have your answer.